12) The change in enthalpy (ΔHorxn) for a reaction is -25.4 kJ/mol . The equilibrium constant for the reaction is 1.4×103 at 298 K. What is the equilibrium constant for the reaction at 602 K ?

13) A reaction has an equilibrium constant of 7.9×103 at 298 K. At 761 K , the equilibrium constant is 0.30. Find ΔHorxn for the reaction.
1
Expert's answer
2020-04-13T02:27:00-0400

The relationship between equilibrium constants at different temperatures, enthaply, and temperatures can be expressed as


"\text{ln}\frac{K_2}{K_1}=\frac{\Delta H}{R}\bigg(\frac{1}{T_1}-\frac{1}{T_2}\bigg),\\
\space\\
K_2=K_1e^{\frac{\Delta H}{R}(\frac{1}{T_1}-\frac{1}{T_2})}=\\ \space\\
=1400\cdot2.72^{\frac{-25400}{8.314}(\frac{1}{298}-\frac{1}{602})}=7.90."

Likewise, we can calculate the change in enthalpy:


"\Delta H=\frac{R\text{ ln}\frac{K_2}{K_1}}{\frac{1}{T_1}-\frac{1}{T_2}}=-41.5\cdot10^3\text{ kJ/mol}."